# Improving the prediction of individual engagement in recommendations using cognitive models

### v1

@@ Abstract.
For public health programs with limited resources, the ability to predict how behaviors change over time and in response to interventions is crucial for deciding when and to whom interventions should be allocated. Using data from a real-world maternal health program, we demonstrate how a cognitive model based on Instance-Based Learning (IBL) Theory can augment existing purely computational approaches. Our findings show that, compared to general time-series forecasters (e.g., LSTMs), IBL models, which reflect human decision-making processes, better predict the dynamics of individuals’ states. Additionally, IBL provides estimates of the volatility in individuals’ states and their sensitivity to interventions, which can improve the efficiency of training of other time series models.
Cognitive Modeling, Instance-based learning, Time-series forecasting, Limited resource allocation, Maternal health, Engagement
@@ 1. Introduction
To reduce dropout, programs can provide interventions designed to increase beneficiary engagement. For example, program staff can make personalized service calls to beneficiaries at risk of dropping out to address concerns about declining participation. However, programs with limited staff and resources cannot intervene on all their beneficiaries. Instead, they must choose a subset of beneficiaries to get the intervention.
@@
Recent approaches to this challenge have adopted the Restless Multi-Armed Bandit (RMAB) resource allocation framework (Killian et al., 2023). Within this framework, the engagement dynamics of each beneficiary (i.e., transition between discrete engagement states) are modeled as a Markov Decision Process (MDP). A centralized planner decides which beneficiaries to intervene upon at each time point based on the dynamics learned from the MDPs. Importantly, beneficiaries can change their engagement state even without an intervention. The most optimal approach in this framework is the Whittle index (Whittle, 1988), which ranks beneficiaries according to the likelihood of becoming and remaining engaged given an intervention.
@@
A key assumption of this framework is that beneficiaries’ dynamics can be modeled as an MDP. However, it is questionable whether these dynamics are indeed Markovian in reality. Recent work demonstrated that these dynamics are often non-Markovian; that is, a beneficiary’s transitions between states (i.e., engaged and disengaged) do depend on their individualized history, and instead developed the Time-series Arm Ranking Index (TARI) which extends RMABs to allow for non-Markovian dynamics and continuous engagement levels (Danassis et al., 2023). According to TARI, each beneficiary is modeled not as a MDP, but using time-series forecasters such as LSTMs (Hochreiter and Schmidhuber, 1997) or transformers (Vaswani et al., 2017). Using these learned time-series models, TARI computes an index that reflects the relative engagement benefit of intervening on a beneficiary versus not intervening at a given time point, and then ranks beneficiaries according to the computed indices.
@@
Within the context described thus far, we propose the use of computational cognitive models of individual beneficiaries instead of LSTMs. Specifically, we will create a model based on a successful account of individual dynamic decision making, Instance-Based Learning Theory (IBLT) (Gonzalez et al., 2003) to represent each beneficiary’s time series activities. Using an IBL model is expected to have at least two potential advantages over general time-series forecasters (TSFs).
@@
First, TSFs need large amounts of data to fine-tune due to the large number of model parameters. Given the relatively sparse amount of data available per beneficiary, existing approaches typically train a single model using data from all available beneficiaries. For instance, Danassis and colleagues (2023) split each beneficiary’s trajectory into a set of fixed-length vectors where each vector represented the intervention history and engagement levels of seven consecutive time steps. Then an LSTM was trained on these vectors, where the task was to predict the engagement level at the next time step (Danassis et al., 2023). In contrast, an IBL cognitive model can be personalized because it can model and predict individual human behavior with less data (Bugbee and Gonzalez, 2022). This is because IBLT makes strong assumptions about the dynamics of human behavior that are derived from prior research on human cognition and decision making. In particular, IBLT accounts for memory effects (i.e., the more recent a data point is, the more it contributes to the prediction) and similarity effects (i.e., engagement levels that are observed under similar contexts should be similar to each other) that general TSFs do not.
@@
Second, the parameters of IBL cognitive models are grounded in psychological constructs such as memory and attention, and are thus more easily interpreted than the parameters of general LSTMs. This advantage in turn enables cognitive models to provide insight about individual differences in cognition that may affect beneficiaries’ behaviors. We demonstrate how these individual cognitive differences can be used to augment existing approaches that rely primarily on general LSTMs.
@@ 1.1. Our Contributions
    Using real-world data from a maternal health program, we show that cognitive models are more accurate compared to LSTMs in capturing the temporal dynamics of individual behaviors and predicting the behaviors of individual beneficiaries.
@@
    We showcase that personalized cognitive models could reveal individual characteristics through weights fitting. We illustrate how to cluster individuals using learnt characteristics.
@@
    We show clustering based on learnt individual characteristics could guide and improve the performance of time series models. Specifically, we show that LSTMs trained in clusters better predict behaviors within those clusters compared to LSTMs trained on random samples of beneficiaries.
@@ 1.2.1. Time-series Arm Ranking Index (TARI)
This training takes place offline. During test time, the model’s parameters are fixed and we use the trained model with iterated multi-step to generate a long-term forecast of future states $s_{t+1},s_{t+2},...,s_{t+H}$ (Taieb et al., 2012). This forecast is then used to calculate the TARI index, which helps us to decide the relative utility of intervening in an arm.
@@
  • Time to disengagement with intervention $u_{n}$: This predicts how long it will take for the arm to become non-engaged if we intervene once at the current timestep (t) and then never intervene again.
  • Time to disengagement without intervention $v_{n}$: This predicts how long it will take for the arm to become non-engaged if we did not intervene at all.
The TARI index for an arm is simply the ratio of these two times. A higher TARI index suggests that intervening now would be more beneficial. Finally, similar to the Whittle index, we choose to act on the $k$ arm with the highest TARI index at each timestep:
@@ 1.2.2. Instance-Based Learning Theory
Instance-Based Learning Theory (IBLT) is a cognitive theory of dynamic decision making, grounded in human learning and memory mechanisms. It has successfully accounted for human behavior in a variety of contexts, ranging from abstract repeated-choice tasks to more real-world search and choice (Nguyen and Gonzalez, 2022) and continuous control tasks (Gonzalez et al., 2003). In particular, in most of these contexts, human participants are required to make a series of decisions while exposed to changing environments. Given IBLT’s previous success in capturing human decision making, we expect that a model based on IBLT would similarly be able to account for the temporal and intervention-related dynamics of beneficiaries’ engagement levels.
At the core of IBLT is the representation of a decision in the form of an instance. Typically, an instance has three parts: the context, the decision itself, and the associated utility. For example, the darkness of the sky would be the context, the choice to bring an umbrella (or not) would be the decision, and the utility may involve the cost of bringing an umbrella added to whether or not you got wet.
IBLT proposes that the agent makes a prediction of the potential utility of a decision based on similar instances stored in memory. For example, it predicts the potential utility of bringing or not bringing an umbrella in situations with similar sky darkness. Decisions are stored as instances in memory and retrieved when the agent is faced with another decision.To predict the utility of each potential action, the past utilities of similar instances are weighted by their probability of retrieval from memory to generate estimates of the expected utility (i.e., a Blended value). The agent then selects the option that has the highest estimated utility. The utilities of instances in memory are updated using the actual outcome observed for a decision made. Thus, the expected initial utility turns into an experienced utility. The updated instances are then bootstrapped to make future decisions.
@@
The contribution of a retrieved instance’s outcome to the final estimated utility of an option depends on the memory activation of the instance. The activation of an instance ($A_{i}$) reflects how readily it comes to mind, which in turn depends on the frequency, recency, and similarity factors according to the following equation as proposed by the ACT-R cognitive architecture (Anderson and Lebiere, 2014).
(1) $A_{i}=ln(\sum(t-t^{\prime})^{-d})+\mu\sum_{j}({w_{j}(Sim(s_{i,j},s_{t,j})-1})+\sigma\xi$
@@
where $d$ is a decay parameter, $\mu$ is a scaling parameter on the overall impact of similarity to activation, $w_{j}$ is the importance (or weight) of an attribute in the similarity calculations, $Sim$ is a similarity function that returns the degree of similarity for a particular attribute $j$ between instance $i$ and the current context, and $\sigma\xi$ is a scaled noise distribution. The activation of an instance is then converted into a retrieval probability as per the Boltzmann softmax function:
@@
where $\tau$ is the temperature parameter. The expected utility value of an option $(a)$ is then computed using all retrieved instances weighted by their retrieval probabilities through blending as proposed in (Gonzalez et al., 2003; Gonzalez and Dutt, 2011):
@@
The agent then chooses the option with the highest blended value $V$.
@@ 2. Proposed Approach
Consistent with IBLT, we account for engagement at the individual level such that each beneficiary is modeled by an IBL model. Each IBL model is trained through model tracing; at each timestep, the model is given an instance that represents the beneficiary’s recent engagement history and current engagement level (see Figure 1a). This instance is then stored in the model’s memory, and is subject to the same temporal and similarity mechanisms that govern its influence on future estimates of the beneficiary’s engagement level.
For IBL models to act as time-series predictors, we use a model’s blended value of the current engagement context as its prediction of the beneficiary’s engagement level. Each IBL instance has two attributes that are derived purely from the beneficiary’s engagement history: the number of time steps since the last intervention and the engagement level on the immediately prior time step. The IBL model determines the similarity between the current context and each instance in memory according to these two attributes, and these similarities are factored into an instance’s activation according to Equation  1.
As mentioned above, the similarity of two instances depends not only on the similarity between attribute values, but also on the weight or influence of the attribute ($w$ in Equation  1). Psychologically, an attribute weight represents how relevant the agent considers an attribute to make similarity judgments between two contexts. To capture potential individual differences between mothers, we personalized an IBL model for a specific mother by finding the combination of attribute weight values that resulted in the smallest training prediction error.
Given the predictions of engagement level, we then follow the TARI algorithm and generate a ranked list of mothers according to the predicted relative benefit in the improvement in engagement each mother receives from an intervention versus not receiving one.
@@ 3.1. Data
We tested our approach on data from a maternal healthcare program operated in partnership with a NGO - ARMMAN (ARMMAN, 2024). The main objective of the program is to improve the health outcomes of expectant and early mothers in vulnerable communities through the dissemination of automated messages about maternal and infancy care. To help increase the engagement of mothers enrolled in the program, ARMMAN staff can deliver a limited number of interventions per week in the form of voice calls in which mothers have the opportunity to ask questions and interact with ARMMAN staff.
@@
The dataset was collected by ARMMAN in 2022 and contains the data of 12,000 mothers over a 40-week period. Two data points were recorded for each mother per week: whether or not she received a call from an ARMMAN staff and the amount of time she spent listening to automated health messages. The latter is used as a proxy for engagement level, normalized to the range [0, 1]. Although the dataset also contained the demographic information of each mother, it was found that there was little or no prediction benefit in incorporating demographic variables as additional features for time-series forecasters (Danassis et al., 2023). Thus, we focus solely on the behavioral variables (presence of intervention and level of engagement).
@@
For an IBL model to be able to predict the effect of an intervention on a mother’s engagement level, it needs to be trained on a trajectory that includes at least one intervention. However, of the 12,000 mothers, only 5,400 mothers received at least 1 intervention early on in their program tenure. We also assume that mothers receive an intervention (or the equivalent of an intervention) upon their enrollment in the program. Furthermore, because the amount of time needed to train all IBL models scales up with the number of mothers, we decided to explore our approach with a subset of 210 mothers out of the 5,400 mothers who received at least one intervention in their actual trajectories.
@@
Figure 1. a) Transforming a beneficiary’s trajectory into IBL instances and LSTM sliding windows. Each IBL instance represents the engagement context at a given timestep and consists of three parts; the engagement level on the immediately prior timestep, the number of timesteps since the last intervention, and the engagement level on the current timestep. In contrast, the LSTM data was constructed by segmenting trajectories into sliding windows of length 7. b) Training / testing setup for comparing IBL and LSTM methods. Different colors indicate data from individual beneficiaries. Under the IBL procedure, one IBL model is trained per beneficiary. Each trained IBL model is then used to make predictions for their respective beneficiaries during testing. Under the LSTM procedure, one LSTM is trained using data from all beneficiaries. The trained LSTM model is then used to make predictions for all beneficiaries during testing. c) LSTM training / testing setups under various IBL-clusters-related conditions. Different colors indicate different cluster types. Note that the LSTM in the Random condition is trained on a random subset of one-third of the training data as compared to the other training conditions.

### v2

@@ Abstract.
For public health programs with limited resources, the ability to predict how behaviors change over time and in response to interventions is crucial for deciding when and to whom interventions should be allocated. Using data from a real-world maternal health program, we demonstrate how a cognitive model based on Instance-Based Learning (IBL) Theory can augment existing purely computational approaches. Our findings show that, compared to general time-series forecasters (e.g., LSTMs), IBL models, which reflect human decision-making processes, better predict how individuals’ behaviors change over time (transition-consistency) and in response to receiving an intervention (intervention-sensitivity). We further show that IBL parameters capture the individual differences in transition-consistency and intervention-sensitivity and that other time series models can use these individual-level IBL parameters to improve their training efficiency.
Limited resource allocation, Restless multi-armed bandit, Time-series forecasting, Cognitive modeling, Instance-based learning
@@ 1. Introduction
To reduce dropout, programs can provide interventions designed to increase beneficiary engagement. For example, program staff can make personalized service calls to beneficiaries at risk of dropping out to address concerns about declining participation. However, programs with limited staff and resources cannot provide interventions to all their beneficiaries. Instead, they must choose a subset of beneficiaries to receive the intervention.
@@
Recent approaches to this challenge have adopted the Restless Multi-Armed Bandit (RMAB) resource allocation framework (Killian et al., 2023). Classically, the RMAB framework models the engagement dynamics of each beneficiary (i.e., the transition between discrete engagement and disengagement states) as a Markov Decision Process (MDP). A centralized planner decides which beneficiaries to intervene on at each time point based on the dynamics learned from the MDPs. Importantly, beneficiaries can change their engagement state even without an intervention. The best-performing approach to this framework is the Whittle index (Whittle, 1988), which ranks beneficiaries by their likelihood of becoming and remaining engaged after getting an intervention.
@@
This framework assumes that beneficiaries’ dynamics—how their level of engagement changes over time—can be modeled as an MDP. However, recent work has demonstrated that these dynamics are often non-Markovian; that is, a beneficiary’s transition between being engaged and disengaged with the program (their state) depends on their individual history (Danassis et al., 2023). As a substitute, Danassis and colleagues (2023) developed the Time-series Arm Ranking Index (TARI), which allows non-Markovian dynamics and continuous engagement levels within the RMAB framework. The TARI approach models beneficiaries using time-series forecasters such as LSTMs (Hochreiter and Schmidhuber, 1997) or transformers (Vaswani et al., 2017) instead of MDPs. Using these learned time-series models, TARI computes an index that captures the relative engagement benefit of intervening on a beneficiary by comparing their expected future engagement with and without an intervention at the current time point and then ranks beneficiaries according to the computed indices.
@@
Within the TARI approach to the RMAB framework, we propose modeling individual beneficiaries with computational cognitive models instead of LSTMs. Specifically, we create a model based on Instance-Based Learning Theory (IBLT) (Gonzalez et al., 2003) to represent each beneficiary’s time series activities. We expect using an IBL model will convey two advantages over general time-series forecasters (TSFs).
@@
First, IBL models are personalized to the individual instead of trained across the entire dataset, allowing the models to capture individual differences in behavioral dynamics (i.e., individual-level differences in the response to interventions and general engagement patterns). Because TSFs need large amounts of data to fine-tune their many parameters and there is relatively sparse data available per beneficiary, existing approaches train a single model using the combined data from all available beneficiaries. For instance, Danassis and colleagues (2023) create a set of fixed-length vectors from each beneficiary’s trajectory. Each vector represents the intervention history and engagement levels of seven consecutive time steps. Then an LSTM trains on these vectors to predict the engagement level at the next time step given the previous seven (Danassis et al., 2023). In contrast, an IBL model has enough cognitively grounded structure that it can be applied to an individual with sparse data (Bugbee and Gonzalez, 2022). This is because IBL theory makes strong assumptions about the cognition of human decision making derived from prior research. In particular, IBL theory accounts for memory effects (i.e., the more recent a data point is, the more it contributes to the prediction) and similarity effects (i.e., engagement levels that are observed under similar contexts should be similar to each other) that general TSFs do not.
@@
Second, the parameters of IBL cognitive models are grounded in psychological constructs such as memory and attention, making them more interpretable than the parameters of general LSTMs. This advantage allows cognitive models to shed light on the individual differences that drive beneficiaries’ behaviors. We demonstrate how these individual differences can be used to augment existing approaches that rely primarily on general LSTMs.
@@ 1.1. Our Contributions
    Using real-world data from a maternal health program, we show that cognitive models more accurately capture the temporal dynamics of individual behaviors and predict the behaviors of individual beneficiaries than LSTMs.
@@
    We showcase that personalized cognitive models could reveal individual characteristics through weights fitting. We illustrate how to cluster individuals using learned characteristics.
@@
    We show clustering based on learned individual characteristics could guide and improve the performance of time series models. Specifically, we show that LSTMs trained in clusters better predict behaviors within those clusters compared to LSTMs trained on random samples of beneficiaries.
@@ 1.2.1. Time-series Arm Ranking Index (TARI)
This training takes place offline. During test time, the model’s parameters are fixed, and we use the trained model with iterated multi-step forecasting to generate a long-term forecast of future states $s_{t+1},s_{t+2},...,s_{t+H}$ (Taieb et al., 2012).
@@
  • Time to disengagement with intervention $u_{n}$: The predicted number of timesteps until the arm becomes non-engaged if we intervene at the current timestep (t) and never again.
  • Time to disengagement without intervention $v_{n}$: The predicted number of timesteps until the arm becomes non-engaged with no interventions at any point.
The TARI index for an arm is the ratio of these two times. A higher TARI index suggests that intervening now would be more beneficial. Finally, similar to the Whittle index, we choose to act on the $k$ arm with the highest TARI index at each timestep:
@@ 1.2.2. Instance-Based Learning Theory
Instance-Based Learning Theory (IBLT) is a cognitive theory of dynamic decision making, grounded in human learning and memory mechanisms. It has successfully accounted for human behavior in a variety of contexts, ranging from abstract repeated-choice tasks to more real-world search and choice (Nguyen and Gonzalez, 2022) and continuous control tasks (Gonzalez et al., 2003). In most of these contexts, human participants are required to make a series of decisions while exposed to changing environments.
The essence of IBLT is storing past decisions in the form of an instance. An instance has three parts: the context, the choice made, and the associated utility. For example, the cloudiness of the sky could be the context, the choice to bring an umbrella could be the choice, and the utility could be the inconvenience of bringing an umbrella when it does not rain (supposing it did not rain in this instance).
IBL agents compute the expected utility of a choice by considering past instances where they made that choice, weighted by the contextual similarity to the present. For example, an IBL agent would compute the expected utility of bringing an umbrella by considering its past experiences with an umbrella, most strongly considering the experiences with the most similar cloudiness to the current moment.
More precisely, the expected value of a choice is the weighted average of the utilities from all the past instances where that choice was made weighted by the retrieval probability (i.e., the probability of remembering that instance) to generate estimates of the choice’s expected utility called a blended value. The agent will deterministically make the available choice with the highest computed blended value (expected utility). The utilities of instances in memory are the actual outcome the agent observed when it chooses, regardless of what it predicted. After choosing, a new instance is added to the agent’s memory, reflecting what happened. If the exact combination of context, choice, and utility has appeared before, that instance instead has its frequency and recency information updated.
@@
The contribution of an instance’s utility to a choice’s expected value depends on the instance’s memory activation (the salience of that instance given the current context). The activation of an instance ($A_{i}$) reflects how readily it comes to mind, which in turn depends on the frequency, recency, and similarity factors according to the following equation as proposed by the ACT-R cognitive architecture (Anderson and Lebiere, 2014).
(1) $A_{i}=ln(\sum_{j}(t-t_{i,j}^{\prime})^{-d})+\mu\sum_{k}({w_{k}(Sim(s_{i,k},s_{t,k})-1})+\sigma\xi$
@@
Three expressions additively determine an instance’s activation (salience in memory). The first expression is the contribution of frequency and recency where $t$ is the current timestep, $t_{i,j}$ is the timestep of the $j$th appearance of instance $i$, and $d$ is a decay parameter. The second expression is the contribution of similarity where $\mu$ is a scaling parameter on the overall impact of similarity to activation, $w_{k}$ is the importance (or weight) of an attribute in the similarity calculations, $Sim$ is a similarity function that returns the degree of similarity for a particular attribute $k$ between instance $i$ ($s_{i,k}$) and the current context ($s_{t,k}$). The third expression adds some randomness to the memory retrieval process where $\sigma\xi$ is a scaled noise distribution. The activation of an instance is then converted into a retrieval probability using the Boltzmann softmax function, where $\tau$ is the temperature parameter:
@@
The IBL model computes the expected utility of a choice $(a)$ from the utility $u_{i}$ of each retrieved instance $i$ weighted by their respective retrieval probability $P_{i}$ (Gonzalez et al., 2003; Gonzalez and Dutt, 2011):
@@
The agent then makes the choice with the highest expected utility, $V(a)$.
@@ 2. Proposed Approach
We model each beneficiary with their own individual-level IBL model. We do not have the IBL model make a choice. Instead, we use the IBL’s expected utility calculation to predict the engagement level in the next period (see Figure 1a). Because we only use the model’s blended value, the choice part of the instance is always held constant. The utility of an instance is a beneficiary’s engagement level (thus, the expected utilities are weighted averages of past engagement levels then used to predict future engagement levels). The context has two components derived from a beneficiary’s history: (1) the engagement level of the previous period and (2) the number of timesteps since the last intervention. The IBL model determines the similarity between the current context and each instance in memory according to these two attributes, and these similarities are factored into an instance’s activation according to Equation  1.
The IBL models are trained through model tracing: recording each timestep of an individual beneficiary as an instance with the relevant context (previous engagement level and time since last intervention) and utility (engagement level). Thus, the model “traces” the beneficiary it is modeling by giving itself memories as if it were the beneficiary that produced the data.
As shown in Equation  1, the similarity of two instances depends not only on the similarity between attribute values, but also on the weight or influence of the attribute ($w$ in Equation  1). Psychologically, an attribute weight represents how relevant the agent considers an attribute to make similarity judgments between two contexts. To capture potential individual differences between beneficiaries, we personalized an IBL model for a specific beneficiary by finding the combination of attribute weight values that resulted in the smallest training prediction error.
Given the predictions of engagement level, we then follow the TARI algorithm and generate a ranked list of beneficiaries according to the predicted relative benefit in the improvement in engagement each beneficiary receives from an intervention versus not receiving one.
@@ 3.1. Data
We tested our approach on data from a maternal healthcare program operated by an NGO called ARMMAN (ARMMAN, 2024). This program sends automated messages about maternal and infant care to expectant and recent mothers in vulnerable communities. However, over the many weeks of pregnancy and postnatal care, many mothers stop engaging with the automated messages. ARMMAN has a limited capacity to intervene: Program staff can call mothers individually, allowing them to hear from a real person and ask questions, hopefully increasing their future engagement.
@@
ARMMAN collected the data in 2022 from 12,000 mothers over 40 weeks. The data have two recorded values per mother per week: the amount of time she spent listening to the automated health message (engagement level, recorded as a number in $[0,1]$) and if she received a call from ARMMAN staff (if she received an intervention, recorded as a binary value). The data also contain mothers’ demographic information; however, previous work found little or no prediction benefit in incorporating demographic variables as additional features for time-series forecasters (Danassis et al., 2023). Thus, we do not consider them in this approach.
@@
For an IBL model to predict an intervention’s effect on a mother’s engagement level, it needs to be trained on a trajectory that includes at least one intervention. However, of the 12,000 mothers, only 5,400 mothers received at least one intervention early on in their program tenure. We also assume that mothers receive an intervention (or the equivalent of an intervention) upon their enrollment in the program. Because the time needed to train all the IBL models scales with the number of mothers, we explored our approach with a subset of 210 mothers out of the 5,400 mothers who received at least one intervention in their actual trajectories.
@@
Figure 1. a) Transforming a beneficiary’s trajectory into IBL instances and LSTM sliding windows. Each IBL instance represents the engagement context at a given timestep and consists of three parts: the engagement level on the immediately prior timestep, the number of timesteps since the last intervention, and the engagement level on the current timestep. In contrast, the LSTM data was constructed by segmenting trajectories into sliding windows of length 7. b) Training / testing setup for comparing IBL and LSTM methods. Different colors indicate data from individual beneficiaries. Under the IBL procedure, one IBL model is trained per beneficiary. Each trained IBL model is then used to predict their respective beneficiaries during testing. Under the LSTM procedure, one LSTM is trained using data from all beneficiaries. The trained LSTM model is then used to make predictions for all beneficiaries during testing. c) LSTM training / testing setups under various IBL-clusters-related conditions. Different colors indicate different cluster types. Note that the LSTM in the Random condition is trained on a random subset of one-third of the training data as compared to the other training conditions.